OpenStudy (anonymous):

find the area of parallelogram ABCD given AB=12cm, AD=20cm and angle A = 30 degrees

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

area=absinA=12*20*sin30=120 cm^2

OpenStudy (anonymous):

can u do that without the sin - i am just in trig - is it the height 6sqrt3

OpenStudy (anonymous):

not in trig sorry just in geometry

OpenStudy (anonymous):

ok, in geometry u use the formula area=bh. but to calculate the height from the given length of sides and the given angle u must use trigonometry.

OpenStudy (anonymous):

well since a 30 60 90 hypotenuse is s which is 12 and height is sqrt3/2 and bottom is s/2

OpenStudy (anonymous):

so figured out height is 6sqrt 3?

OpenStudy (anonymous):

if u take the 12 cm side as a base, the height is 20sin30=10. then the area becomes 12*10=120. if u take the 20 cm side as a base, the height is 12sin30=6. then the area becomes 20*6=120. thus the length of the height depends on which side u take as a base.

OpenStudy (anonymous):

i need it without trig i am in geometry - i know that it is a 30 60 90 triangle

OpenStudy (anonymous):

yes, but the possible heights are 10 and 6, but not 6sqrt3.

OpenStudy (ranga):

In a 30-60-90 triangle, the corresponding sides are: 1:sqrt(3):2 so the side opposite 30 degrees is 1/2 the length of the hypotenuse. So the height of the parallelogram is either 20/2 or 12/2. Area = base * height = 12 * 20/2 = 120 or 20 * 12/2 = 120.